OpenStudy (abbi):

Solve the system of equations x + 3y = 32 -3x + 2y = 3

OpenStudy (anonymous):

from the first equation we have: x = 32 - 3y Substituting this value of x into the second equation we find: -3(32 - 3y) + 2y = 3 -96 + 9y + 2y = 3 11y = 99 y = ? With the value of Y we can find X with: x = 32 - 3( ? ) Did yu understand?

OpenStudy (abbi):

if x = 29 and y = 3 the the first equation would = 38 then and not 32 correct

OpenStudy (anonymous):

are you sure y=3? (:

OpenStudy (abbi):

no im think 9 b/c 11y=99 would be 11*9=99

OpenStudy (anonymous):

yup, y = 9